OpenStudy (anonymous):

√2x+8 - √x+11 = -1 find the solution set

OpenStudy (adamaero):

Is the 2 and x the only ones in sq root?

OpenStudy (anonymous):

(2x+*) is one sqrt and (x+11) is another one :)

OpenStudy (adamaero):

ok, I just about got it

OpenStudy (anonymous):

okay i go the answer as 0 but apparently is incorrect

OpenStudy (anonymous):

my guess is 14 using no algebra

OpenStudy (adamaero):

so you rearranged sides and went from there?

OpenStudy (anonymous):

yeah, i squared both sides and used FOIL but idk

OpenStudy (anonymous):

to hell with foil

OpenStudy (anonymous):

lets be smart

OpenStudy (anonymous):

agreed but thats how i was taught D:

OpenStudy (anonymous):

inside the radicals are \(2x+8\) and \(x+11\) right?

OpenStudy (anonymous):

yesssss

OpenStudy (anonymous):

ok lets focus on \(x+11\) in order for it to be a perfect square like say \(16\) or \(25\) \(x\) what could \(x\) be?

OpenStudy (anonymous):

and the answer is, it could be \(5\) since \(5+11=16\) BUT \(2\times 5+8=18\) is NOT a perfect square, so we toss that out

OpenStudy (anonymous):

x=14? idk if thats what you mean

OpenStudy (anonymous):

so what if \(x+11=25\)? then \(x=14\) and \[2\times 14+8=36\] another perfect square !! got it on the second try

OpenStudy (anonymous):

the check is easy \[\sqrt{2\times 14+8}-\sqrt{14+11}=\sqrt{36}-\sqrt{25}=6-5=1\]

OpenStudy (anonymous):

ooh snap !! you want \(-1\) right !!

OpenStudy (anonymous):

superb. then the solution is 14! thank you, i always mess up with the foil method D;

OpenStudy (anonymous):

no no i was wrong!!

OpenStudy (anonymous):

wait... why D: it looks right?

OpenStudy (anonymous):

you want \(-1\) not \(1\) i read it wrong

OpenStudy (anonymous):

oh okay. so what would i do then?

OpenStudy (anonymous):

so lets make \(x+11=9\) another perfect square

OpenStudy (anonymous):

that makes \(x=-2\) and it makes \[2x+8=2\times (-2)+8=4\] that's better

OpenStudy (anonymous):

now we get it \[\sqrt{2\times -1+8}-\sqrt{-2+11}=\sqrt4-\sqrt9=2-3=-1\]

OpenStudy (anonymous):

it is always better to think than it is to use algebra algebra is for math teachers, thinking is for humans but if you want to do it the math teacher way we can do that too

OpenStudy (anonymous):

would there be a solution set?

OpenStudy (anonymous):

more then just 8?

OpenStudy (anonymous):

yes, it is \(\{-2\}\)

OpenStudy (anonymous):

okay, now can you explain to me how you got to find that answer so quickly?

OpenStudy (anonymous):

yes, that is what i was trying to do let me say it in plain english

OpenStudy (anonymous):

you have two radicals inside one radical is \(2x+8\) and inside the other is \(x+11\)

OpenStudy (anonymous):

in order for the answer to come out to a nice whole number, the number inside BOTH radicals has to be a perfect square, like \[1,4,9,16,25,36,49,64,...\]

OpenStudy (anonymous):

how likely is that? not very the problem has to be cooked up carefully to make it work so i made a guess i guessed that \(x+11=16\) making \(x=5\) but that did not make \(2x+8\) a perfect square, it made it 18 toss that out

OpenStudy (anonymous):

then i tried \(x+11=25\) making \(x=14\) and was happy because \(2x+8=36\) but i was again mistaken because that gave \(6-5=1\) and you wanted \(-1\) too big try \(x+11=9\) which makes \(x=-2\) and lo and behold \(2x+8=4\) done

OpenStudy (anonymous):

now we can do it the donkey way if you like start with \[\sqrt{2x+8}-\sqrt{x+11}=-1\] then add \(\sqrt{x+11}\)to both sides giving \[\sqrt{2x+8}=\sqrt{x+11}-1\] then square both sides \[2x+8=x+11-2\sqrt{x+11}+1\] ick

OpenStudy (anonymous):

now we have to put all the stuff not in the racial on the left and square again!!

OpenStudy (anonymous):

plus it is easy to make a mistake when squaring \((\sqrt{x+11}-1)^2\)

OpenStudy (anonymous):

okay! im going to take note of this! thank you so much!!!! :D

OpenStudy (anonymous):

we can do it if you like, we will still get \(x=-2\)

OpenStudy (anonymous):

and if you cannot guess correctly i spose you have to do it the math teacher way, but try to make a good guess first especially if you have two radicals yw